OpenStudy (anonymous):

Solving trig equation on the interval 0

OpenStudy (anonymous):

\[3-3\sin \theta =6\cos^2\theta\]

OpenStudy (anonymous):

I solved it to equal \[0=6\sin^2\theta+3\sin \theta-2\] just need help factoring

OpenStudy (astrophysics):

Let sin theta = x, and solve as you would regularly than plug it back in after :)

OpenStudy (anonymous):

That's what I'm having problems with, the factoring part. Whether it's sin theta or x doesn't make a difference to me. I have a lot of problems factoring :(

OpenStudy (mertsj):

\[3-3\sin \theta =6(1-\sin ^2\theta)\] \[3-3\sin \theta =6-6\sin ^2\theta \] \[6\sin ^2\theta -3\sin \theta -3=0\]\[2\sin ^2\theta-\sin \theta-1=0\]

OpenStudy (mertsj):

\[(2\sin \theta+1)(\sin \theta-1)=0\]

OpenStudy (mertsj):

Set each factor equal to 0 and solve.

OpenStudy (anonymous):

ah I did \[3-3\sin \theta = 1+6\sin^2\theta\] then got \[0=6\sin^2\theta+3\sin \theta -2\] I didnt realize you were supposed to write it out like that which made the rest of it get messed up. Thanks.